OpenStudy (anonymous):

Solve the differential equation :- (y+3x^2)dy/dx=x

OpenStudy (anonymous):

y=ux2 dy=x2du+2xudx

OpenStudy (anonymous):

You sure? :/ So you too y as ux^2, eh?

OpenStudy (anonymous):

actually u=y/x2 is a two variable function to solve the de

OpenStudy (anonymous):

Hm, that's perfectly fine.. It's just, i don't get the answer :(

OpenStudy (anonymous):

ok just put the new variable u instead y then you reach the answer.

OpenStudy (anonymous):

(ux2+3x2)(x2du+2xudx)=xdx

OpenStudy (anonymous):

(u+3)x4du+2(u+3)x3udx=xdx